# Louis Serrano1  Armand Kassaï Koupaï1  Thomas X Wang1 Pierre Erbacher 2Patrick Gallinari1,3 1Sorbonne Université, CNRS, ISIR, 75005 Paris, France

### v2

@@ Abstract
Solving time-dependent parametric partial differential equations (PDEs) is challenging, as models must adapt to variations in parameters such as coefficients, forcing terms, and boundary conditions. Data-driven neural solvers either train on data sampled from the PDE parameters distribution in the hope that the model generalizes to new instances or rely on gradient-based adaptation and meta-learning to implicitly encode the dynamics from observations. This often comes with increased inference complexity. Inspired by the in-context learning capabilities of large language models (LLMs), we introduce Zebra, a novel generative auto-regressive transformer designed to solve parametric PDEs without requiring gradient adaptation at inference. By leveraging in-context information during both pre-training and inference, Zebra dynamically adapts to new tasks by conditioning on input sequences that incorporate context trajectories or preceding states. This approach enables Zebra to flexibly handle arbitrarily sized context inputs and supports uncertainty quantification through the sampling of multiple solution trajectories. We evaluate Zebra across a variety of challenging PDE scenarios, demonstrating its adaptability, robustness, and superior performance compared to existing approaches. GitHub page: https://github.com/LouisSerrano/zebra.
@@ 1 Introduction
Training partial differential equation (PDE) solvers is a challenging task due to the variety of behaviors that can arise in physical phenomena, and neural solvers have limited generalization capability(Chen et al., 2018; Raissi et al., 2019; Li et al., 2021). We tackle the parametric PDE problem (Cohen & Devore, 2015), where a model is trained on trajectories defined by varying PDE parameters with the goal of generalizing across a wide range of parameters. The parameters may include initial and boundary conditions, physical coefficients, and forcing terms. We focus on pure data-driven approaches that do not leverage any prior knowledge on the underlying equations.
A natural approach to this problem is to sample from the parameter distribution, i.e., to train using different PDE instances or parameter values, along with multiple trajectories for each PDE instance. This requires a training set representative of the distribution of the underlying dynamical system, which is difficult to meet in practice given the complexity of physical phenomena. Other approaches explicitly condition on specific PDE parameters, (Brandstetter et al., 2022b; Takamoto et al., 2023) relying on the availability of such prior knowledge. This requires a physical model of the observed system, making the incorporation of PDE parameters into neural solvers challenging beyond basic PDE coefficients. An alternative approach involves online adaptation to new PDE instances by leveraging observations from novel environments. Here we consider that an environment is charaterized by a set of parameters. This adaptation is often implemented through meta-learning, where the model is trained on a variety of simulations corresponding to different environments—i.e., varying PDE parameter values—so that it can quickly adapt to new, unseen PDE simulation instances using a few trajectory examples(Kirchmeyer et al., 2022; Yin et al., 2022). This method offers a high flexibility but requires gradient updates for adaptation, adding computational overhead. Another common setting involves leveraging historical data to condition the neural network, allowing it to generalize to new PDE instances without retraining (Li et al., 2021; McCabe et al., 2023). Again the generalization ability is limited to dynamics close to the ones used for training. Exploring another direction and motivated by the successes encountered in natural language processing and vision, some authors have begun investigating the development of foundation models for spatio-temporal dynamic physical processes (Subramanian et al., 2023; Herde et al., 2024; McCabe et al., 2023). This approach involves training a large model on a variety of physics-based numerical simulations with the expectation that it will generalize to new situations or equations. While they consider multiple physics we focus on solving parametric PDEs, i.e. multiple variations of the same physical phenomenon.
We explore here a new direction inspired by the successes of in-context learning (ICL) and its ability to generalize to downstream tasks without retraining (Brown et al., 2020; Touvron et al., 2023). We propose a framework, denoted Zebra, relying on in-context pretraining (ICP), for solving parametric PDEs and learning to condition neural solvers to adapt fast to new situations or said otherwise for solving for new parameter values. As for ICL in language the model is trained to generate appropriate responses given context examples and a query. The context examples could be trajectories from the same dynamics starting from different initial conditions, or simply a brief history of past system states for the target trajectory. The query will consist for example of an initial state condition, that will serve as inference starting point for the forecast. This approach offers key advantages compared to existing methods. It can leverage contexts of different types and sizes, it requires only a few context examples to adapt to new dynamics and can can handle as well 0-shot learning. It allows us to cover a large variety of situations.
@@
On the technical side, Zebra introduces a novel generative autoregressive solver for parametric PDEs. It employs an encode-generate-decode framework: first, a vector-quantized variational auto-encoder (VQ-VAE) (Oord et al., 2017) is learnt to compress physical states into discrete tokens and to decode it back to the original physical space. Next, a generative autoregressive transformer is pre-trained using a next token objective. To leverage the in-context properties of the model, Zebra is directly pretrained on arbitrary-sized contexts such as extra trajectories or historical states of the target dynamics. At inference, Zebra can handle varying context sizes for conditioning and support uncertainty quantification, enabling generalization to unseen PDE parameters without gradient updates.
@@
  • We introduce a generative autoregressive transformer for modeling physical dynamics. It operates on compact discretized representations of physical state observations. This discretization is performed through a VQ-VAE. The encoder tokenizes observations into sequences of tokens, while the decoder reconstructs the original states. This framework represents the first successful application of generative modeling using quantized representations of physical systems.
  • To harness the in-context learning strengths of autoregressive transformers, we develop a new pretraining strategy that conditions the model on historical states or example trajectories with similar dynamics, allowing it to handle arbitrary-sized context token inputs.
  • We evaluate Zebra on a range of parametric PDEs on two distinct settings. In the first, the model infers dynamics from a context trajectory that shares similar behavior with the target but differs in initial conditions, representing a one-shot setting. Zebra’s performance is benchmarked against domain-adaptation baselines specifically trained for such tasks. In the second scenario, only a limited number of historical frames of the target trajectory are available, requiring the model to deduce the underlying dynamics solely from these inputs. Zebra consistently demonstrates competitive performance across both evaluation contexts.
@@ 2 Problem setting
2.1 Solving parametric PDE
We aim to solve parametric time-dependent PDEs beyond the typical variation in initial conditions. Our goal is to train models capable of generalizing across a wide range of PDE parameters. To this end, we consider time-dependent PDEs with different initial conditions, and with additional degrees of freedom, namely: (1) coefficient parameters — such as fluid viscosity or advection speed — denoted by vector $\mu$ ; (2) boundary conditions $\mathcal{B}$, e.g. Neumann or Dirichlet; (3) forcing terms $\delta$, including damping parameter or sinusoidal forcing with different frequencies. To simplify notation we denote $\xi:=\{\mu,\mathcal{B},\delta\}$ and we define $\mathcal{F}_{\xi}$ as the set of PDE solutions corresponding to the PDE parameters $\mu$, boundary conditions $\mathcal{B}$ and forcing term $\delta$, and refer to $\mathcal{F}_{\xi}$ as a PDE partition. Formally, a solution ${\bm{u}}(x,t)$ within $\mathcal{F}_{\xi}$ satisfies:
$\frac{\partial{\bm{u}}}{\partial t}$ $=F\left(\delta,\mu,t,x,{\bm{u}},\frac{\partial{\bm{u}}}{\partial x},\frac{\partial^{2}{\bm{u}}}{\partial x^{2}},\ldots\right),\quad\forall x\in\Omega,\forall t\in(0,T]$ (1)
$\mathcal{B}({\bm{u}})(x,t)$ $=0,\quad\forall x\in\partial\Omega,\forall t\in(0,T]$ (2)
${\bm{u}}(0,x)$ $={\bm{u}}^{0},\quad\forall x\in\Omega$ (3)
@@ 2.1 Solving parametric PDE
Figure 1: Zebra Framework for solving parametric PDEs. 1) A finite vocabulary of physical phenomena is learned by training a VQ-VAE on spatial representations. 2) During the pretraining, multiple trajectories sharing the same dynamics are tokenized and concatenated into a common sequence $S$. A transformer is used to predict the next tokens in these sequences, conditioned on the context. This enables the model to perform both zero-shot and few-shot generation, without gradient-based updates.
2.2 Generalization for parametric PDE
@@
Solving time-dependent parametric PDEs requires developing neural solvers capable of generalizing to a whole distribution of PDE parameters. In practice, changes in the PDE parameters often lead to distribution shifts in the trajectories which makes the problem challenging. Different directions are currently being explored briefly reviewed below. We focus on pure data-driven approaches that do not make use of any prior knowledge on the equations. We make the assumption that the models are learned from numerical simulations so that it is possible to generate from multiple parameters. This emulates real situations where for example, a physical phenomenon is observed in different contexts.
0-shot learning with temporal conditioning
A first direction consists in adapting the classical ERM framework to parametric PDE solving by sampling multiple instances of a PDE, in the hope that this will generalize to unseen conditions in a 0-shot setting. It is usually assumed that for both learning and inference, a sequence of past states is provided as initial input to the model, leveraging its potential to infer the dynamics characteristics in order to forecast future values. The neural solver $\mathcal{G}_{\theta}$ is then conditioned by a sequence of past states for a trajectory ${\bm{u}}^{t-m\Delta t:t}:=({\bm{u}}^{t-m\Delta t},\ldots,{\bm{u}}^{t})$ where $m\geq 1$. Depending on the architecture, this can be implemented by stacking the information in the channel dimension (Li et al., 2021), or by creating an additional temporal dimension as done in video prediction contexts (McCabe et al., 2023; Ho et al., 2022). This approach makes an implicit i.i.d. assumption on the training - test distributions which is often not met with dynamical phenomena. It offers a limited flexibility in cases where only limited historical context is accessible.
Few-shot learning by fine tuning
Another category of methods leverages fine tuning. As for the 0-shot setting above, a model is pretrained on a distribution of the PDE parameters. At inference, for a new environment, fine tuning is performed on a sample of the environment trajectories. This approach often relies on large fine tuning samples and involves updating all or a subset of parameters (Subramanian et al., 2023; Herde et al., 2024).
Adaptive conditioning
@@
A more flexible approach relies on adaptation at inference time through meta-learning. It posits that a set of environments $e$ are available from which trajectories are sampled, each environment $e$ being defined by specific PDE parameter values (Zintgraf et al., 2019a; Kirchmeyer et al., 2022). The model is trained from a sampling from the environments distribution to adapt fast to a new environment. The usual formulation is to learn shared and specific environment parameters $\mathcal{G}_{\theta+\Delta\theta_{\xi}}$, where $\theta$ and $\Delta\theta_{\xi}$ are respectively the shared and specific parameters. At inference, for a new environment, only a small number of parameters $\theta_{\xi}$ is adapted from a small sample of observations.
Table 1: Key distinctions with Baselines. Zebra is the only method that supports both adaptive conditioning, temporal conditioning, and does not require gradient computations at inference.
Method Adaptive conditioning Temporal conditioning In-context
CAPE
CODA
MPP
Zebra
@@ 3 Zebra Framework
We introduce Zebra, a novel framework designed to solve parametric PDEs through in-context learning and flexible conditioning. Zebra utilizes an autoregressive transformer to model partial differential equations (PDEs) within a compact, discrete latent space. A spatial CNN encoder is employed to map physical spatial observations into these latent representations, while a CNN decoder accurately reconstructs them. As illustrated in Figure 1, our pretraining pipeline consists of two key stages: 1) Learning a finite vocabulary of physical phenomena, and 2) Training the transformer using an in-context pretraining strategy, enabling the model to effectively condition on contextual information. At inference, Zebra allows both adaptive and temporal conditioning through in-context learning (Table 1).
@@ 3.1 Learning a finite vocabulary of physical phenomena
In order to leverage the auto-regressive transformer architecture and adopt a next-token generative pretraining, we need to convert physical observations into discrete representations. To keep the modeling with the transformer computationnaly tractable, we do not quantize the observations directly but rather quantize compressed latent representations by employing a VQVAE (Oord et al., 2017).
Our encoder spatially compresses the input function ${\bm{u}}^{t}$ by reducing its spatial resolution $H\times W$ to a lower resolution $h\times w$ while increasing the channel dimension to $d$. This is achieved through a convolutional model $\mathcal{E}_{w}$, which maps the input to a continuous latent variable $\mathbf{z}^{t}=\mathcal{E}_{w}({\bm{u}}^{t})$, where $\mathbf{z}^{t}\in\mathbb{R}^{h\times w\times\times d}$. The latent variables are then quantized to discrete codes $\mathbf{z}_{q}^{t}$ using a codebook $\mathcal{Z}$ of size $K=|\mathcal{Z}|$ and through the quantization step $q$. For each spatial code $\mathbf{z}^{t}_{[ij]}$, the nearest codebook entry $z_{k}$ is selected:
The training objective is:
where the first term is the Relative L2 loss commonly used in PDE modeling, and the second term is the commitment loss, ensuring encoder outputs are close to the codebook entries. The parameter $\alpha$, set to 0.25, balances the two components. We provide additional details on the architecutre in Appendix C.
Once this training step is done, we can tokenize a trajectory ${\bm{u}}^{t:t+m\Delta t}$ by applying our encoder in parallel on each timestamp to obtain discrete codes $\mathbf{z}_{q}^{t:t+m\Delta t}$ and retrieve the corresponding index entries $s^{t:t+m\Delta t}$ from the codebook $\mathcal{Z}$. Similarly, we detokenize discrete indices with the decoder.

### v3

# Zebra: In-Context Generative Pretraining for Solving Parametric PDEs

@@ Abstract
Solving time-dependent parametric partial differential equations (PDEs) is challenging for data-driven methods, as these models must adapt to variations in parameters such as coefficients, forcing terms, and initial conditions. State-of-the-art neural surrogates perform adaptation through gradient-based optimization and meta-learning to implicitly encode the variety of dynamics from observations. This often comes with increased inference complexity. Inspired by the in-context learning capabilities of large language models (LLMs), we introduce Zebra, a novel generative auto-regressive transformer designed to solve parametric PDEs without requiring gradient adaptation at inference. By leveraging in-context information during both pre-training and inference, Zebra dynamically adapts to new tasks by conditioning on input sequences that incorporate context example trajectories. As a generative model, Zebra can be used to generate new trajectories and allows quantifying the uncertainty of the predictions. We evaluate Zebra across a variety of challenging PDE scenarios, demonstrating its adaptability, robustness, and superior performance compared to existing approaches.
Machine Learning, PDE, Neural Surrogates, ICML
@@ 1 Introduction
A major challenge for training neural solvers for time dependent partial differential equations (PDEs) or more generally for modeling spatio-temporal dynamics is to capture the variety of behaviors arising from complex physical phenomena. In particular, neural solvers, trained from a limited number of situations often fail to generalize to new physical contexts and situations (Chen et al., 2018; Raissi et al., 2019; Li et al., 2021; Koupaï et al., 2024).
We address the parametric PDE problem (Cohen & Devore, 2015), where the goal is to train models on a limited set of scenarios representing a given physical phenomenon so that they can generalize across a wide range of new contexts, including different PDE parameters. These parameters may encompass initial and boundary conditions, physical coefficients, and forcing terms. In this work, we focus on purely data-driven approaches that do not incorporate prior knowledge of the underlying equations.
A basic approach to this problem is to sample from the distribution of physical parameters, i.e., to train on different instances of a PDE characterized by varying parameter values, with the goal of generalizing to unseen instances. This approach relies on an i.i.d. assumption and requires a training set that adequately represents the distribution of the underlying dynamical system—a condition that is often difficult to satisfy in practice due to the complexity of physical phenomena. Other approaches explicitly condition on specific PDE parameters (Brandstetter et al., 2022b; Takamoto et al., 2023), relying on the availability of such prior knowledge. This assumes that a physical model of the observed system is known, making the incorporation of PDE parameters into neural solvers challenging beyond physical quantities. Moreover, in many cases, this prior knowledge is incomplete or entirely unknown.
An alternative approach involves adaptation to new PDE instances by leveraging observations from novel environments. Here we consider that an environment is characterized by a set of parameters. For data-driven models, adaptation is often performed through fine tuning, which usually requires a significant amount of examples for the new environment. This is for example the setting adopted in many recent development of foundation models (Subramanian et al., 2023; Herde et al., 2024; McCabe et al., 2023). This involves training a large model on a variety of physics-based numerical simulations, each requiring a large amount of simulations with the expectation that it will generalize.
More principled frameworks for adaptation leverage meta-learning, where the model is trained on simulations corresponding to different environments—i.e., varying PDE parameter values—so that it can quickly adapt to new and unseen PDE simulation instances using a few trajectory examples (Park et al., 2023; Kirchmeyer et al., 2022; Yin et al., 2022). These flexible methods rely on gradient updates for adaptation, adding computational overhead.
We explore a new direction for adaptation inspired by the successes of in-context learning (ICL) in natural language processing (NLP) and its demonstrated ability to generalize to downstream tasks without retraining (Brown et al., 2020; Touvron et al., 2023). We propose a framework, denoted Zebra, relying on ICL for solving parametric PDEs with new parameter values, without any additional update of the model parameters.
As for ICL in NLP, the model is trained to generate appropriate responses given context examples and a query. The context examples will be trajectories from the same dynamics starting from different initial conditions. The query will consist for example of a new initial state condition, that will serve as inference starting point for the new forecast. The proposed model is inspired from NLP approaches: it is a causal generative model that processes discrete token sequences encoding observations. It is trained to model the trajectory distributions of parameteric PDEs. This approach offers key advantages and greater flexibility compared to existing methods. It can leverage contexts of different types and sizes, requires only a few context examples to adapt to new dynamics, and allows us to cover a wide range of situations. It provides enhanced capabilities compared to more classical deterministic forecasting models. Notably, generative probabilistic models have been developed for physical problems such as weather forecasting (Price et al., 2023; Couairon et al., 2024) and even PDE solving (Lippe et al., 2023), demonstrating superior performance and capabilities over their deterministic counterparts. However, their setting is different, as they rely on diffusion models and are neither designed for adaptation nor intended to address the parametric PDE problem.
Some recent works have also explored adaptation through in-context learning for dynamics modeling. The closest to ours is probably (Yang et al., 2023), which also targets adaptation to multiple environments of an underlying physical dynamics through prompting with examples. Their model employs a specific deterministic encoder-decoder architecture and is limited to 1D ODEs or sparse 2D data due to scalability issues. More details and further references are provided in Appendix A.
@@
On the technical side, Zebra introduces a novel generative autoregressive solver for parametric PDEs. It employs an encode-generate-decode framework: first, a vector-quantized variational auto-encoder (VQ-VAE) (Oord et al., 2017) is learned to compress physical states into discrete tokens and to decode it back to the original physical space. Next, a generative autoregressive transformer is pre-trained with arbitrary size context examples of trajectories using a next token objective. At inference, Zebra can handle varying context sizes for conditioning and supports uncertainty quantification, enabling generalization to unseen PDE parameters without gradient updates.
@@
  • We introduce a generative autoregressive transformer for modeling physical dynamics. It operates on compact discretized representations of physical state observations. This framework represents the first successful application of causal generative modeling using quantized representations of physical systems.
  • To harness the in-context learning strengths of autoregressive transformers, we develop a new pretraining strategy that conditions the model on example trajectories with similar underlying dynamics but different initial conditions.
  • Our generative model predicts trajectory distributions. This provides a richer information than deterministic auto-regressive models. This comes with enhanced capabilities including more accurate predictions, uncertainty measures, or the ability to sample and generate new trajectories conditioned on some examples.
  • We propose an accelerated inference procedure that is orders of magnitude faster than existing adaptation methods.
  • We evaluate Zebra in a one-shot adaptation setting, where it must infer the dynamics from a single context trajectory and unroll the dynamics from a new query initial condition (i.e., a single snapshot). The predicted trajectory is compared to a target trajectory governed by the same underlying dynamics as the context but starting from the query initial condition. Zebra’s performance is benchmarked against domain adaptation baselines specifically trained for this task, and it consistently outperforms gradient-based adaptation methods on 2D datasets.
@@ 2 Problem setting
2.1 Solving parametric PDEs
We aim to solve parametric time-dependent PDEs beyond the typical variation in initial conditions. Our goal is to train models capable of generalizing across a wide range of PDE parameters. To this end, we consider time-dependent PDEs with different initial conditions, and with additional degrees of freedom, namely: (1) coefficient parameters — such as fluid viscosity or advection speed — denoted by vector $\mu$ ; (2) boundary conditions $\mathcal{B}$, e.g. Neumann or Dirichlet; (3) forcing terms $\delta$, including damping parameter or sinusoidal forcing with different frequencies. We denote $\xi:=\{\mu,\mathcal{B},\delta\}$ and we define $\mathcal{F}_{\xi}$ as the set of PDE solutions corresponding to the PDE parameters $\mu$, boundary conditions $\mathcal{B}$ and forcing term $\delta$, and refer to $\mathcal{F}_{\xi}$ as a PDE environment. Formally, a solution ${\bm{u}}(x,t)$ within $\mathcal{F}_{\xi}$ satisfies:
$\frac{\partial{\bm{u}}}{\partial t}$ $=F\big{(}\delta,\mu,t,x,{\bm{u}},\frac{\partial{\bm{u}}}{\partial x},\frac{\partial^{2}{\bm{u}}}{\partial x^{2}},\dots\big{)},$ (1)
$\quad\quad\quad\forall x\in\Omega,\forall t\in(0,T]$
$\mathcal{B}({\bm{u}})(x,t)$ $=0,\quad\forall x\in\partial\Omega,\forall t\in(0,T]$
${\bm{u}}(0,x)$ $={\bm{u}}^{0},\quad\forall x\in\Omega$
@@ 2.1 Solving parametric PDEs
Figure 1: Zebra’s inference pipeline from context trajectories. The model is provided with a set of example trajectories ($u_{1}$ and $u_{2}$) that share the same underlying dynamics (e.g., identical PDE parameters) but differ in initial conditions. Given a new query initial condition $u_{*}^{0}$, Zebra aims to generate the corresponding trajectory that follows the same physical behavior. Each trajectory and initial condition is first tokenized into flattened index sequences $s_{1}$, $s_{2}$, and $s_{*}^{0}$, which are concatenated using a specific formatting scheme. The transformer then autoregressively generates the remaining tokens to obtain a plausible prediction. Finally, the generated indices are detokenized to reconstruct the trajectory in physical space.
2.2 Adaptation for parametric PDE
@@
Solving time-dependent parametric PDEs requires developing neural solvers capable of generalizing to a whole distribution of PDE parameters. In practice, changes in the PDE parameters often lead to distribution shifts in the trajectories which makes the problem challenging. Different directions are currently being explored and are briefly reviewed below. We focus on pure data-driven approaches that do not make use of any prior knowledge on the equations. We make the assumption that the models are learned from numerical simulations so that it is possible to generate from multiple parameters. This emulates real situations where for example, a physical phenomenon is observed in different contexts.
Fine tuning pre-trained models
The classical strategy for adapting to new settings is to fine tune models that have been pretrained on a distribution of the PDE parameters. This approach often relies on large fine tuning samples and involves updating all or a subset of parameters (Subramanian et al., 2023; Herde et al., 2024). We do not consider this option that has been shown to underperform SOTA adaptation approaches (Koupaï et al., 2024).
Gradient-based adaptation
@@
A more flexible approach relies on adaptation at inference time through meta-learning. It posits that a set of environments $e$ are available from which trajectories are sampled, each environment $e$ being defined by specific PDE parameter values (Zintgraf et al., 2019a; Kirchmeyer et al., 2022). The model is trained from a sampling from the environments distribution to adapt fast to a new environment. The usual formulation is to learn shared and specific environment parameters $\mathcal{G}_{\theta+\theta_{\xi}}$, where $\theta$ and $\theta_{\xi}$ are respectively the shared and specific parameters. At inference, for a new environment, only a small number of parameters $\theta_{\xi}$ is adapted from a small sample of observations. This family of method will be our reference baseline in the following.
@@ 3 Zebra Framework
We introduce Zebra, a novel framework designed to solve parametric PDEs through in-context learning and flexible conditioning. Zebra utilizes a generative autoregressive transformer to model partial differential equations (PDEs) within a compact, discrete latent space. A spatial CNN encoder is employed to map physical spatial observations into these latent representations, while a CNN decoder accurately reconstructs them. Our pretraining pipeline consists of two key stages: 1) Learning a finite vocabulary of physical phenomena, and 2) Training the transformer using an in-context pretraining strategy, enabling the model to effectively condition on contextual information. At inference, Zebra allows to perform in-context learning from context trajectories as illustrated in Figure 1.
@@ 3.1 Learning a finite vocabulary of physical phenomena
In order to leverage the auto-regressive transformer architecture and adopt a next-token generative pretraining, we need to convert physical observations into discrete representations. We do not quantize the observations directly but rather quantize compressed latent representations by employing a VQVAE (Oord et al., 2017), an encoder-decoder architecture with a quantizer component. Our encoder spatially compresses the input function ${\bm{u}}^{t}$ through a convolutional model $\mathcal{E}_{w}$, which maps the input to a continuous latent variable $\mathbf{z}^{t}=\mathcal{E}_{w}({\bm{u}}^{t})$. The latent variables are then quantized to a vector of discrete codes $\mathbf{z}_{q}^{t}$ using a codebook $\mathcal{Z}$ of size $K=|\mathcal{Z}|$ through the quantization component $q$. For each spatial code $\mathbf{z}^{t}_{[ij]}$ in $\mathbf{z}_{q}^{t}$, the nearest codebook entry $z_{k}$ is selected. The decoder $\mathcal{D}_{\psi}$ reconstructs the signal $\hat{{\bm{u}}}^{t}$ from the quantized latent codes $\hat{\mathbf{z}}_{q}^{t}$. Both models are jointly trained to minimize the reconstruction error between the function ${\bm{u}}^{t}$ and its reconstruction $\hat{{\bm{u}}}^{t}=\mathcal{D}_{\psi}\circ q\circ\mathcal{E}_{w}({\bm{u}}^{t})$.
Once this training step is done, we can tokenize a trajectory ${\bm{u}}^{t:t+m\Delta t}$ by applying our encoder in parallel on each timestamp to obtain vectors of discrete codes $\mathbf{z}_{q}^{t:t+m\Delta t}$ and retrieve the corresponding index entries $s^{t:t+m\Delta t}$ from the codebook $\mathcal{Z}$. Similarly, we detokenize discrete indices with the decoder. We provide a brief description of the VQVAE model and details on its architecture in Appendix C.